# all_papers_200910.bib

### v2

@@ 5.3 Examining performance across data sets, outputs, and sample sizes
The substantial uncertainty within model families and individual models demonstrates the importance of investigating the contextual factors’ impacts on model performance. Therefore, we further decompose the predictive performance along the contextual dimensions. Figure 5.3 visualizes the distributions of prediction accuracy in three data sets, and Figure 5.3 examines the prediction accuracy regarding data sets, outputs, and sample sizes. Table 5 reports the model ranking and performance in three panels and nine sub-tables. Within each sub-table, we report the top-10 models with the DCMs highlighted below, conditioning on specific data sets, outputs, and sample sizes. With such analysis, we could examine the performance consistency of individual models and model families across contexts.
Fig. 4.: Prediction accuracy by data sets (dashed and dotted lines represent mean and median values)
(a) NHTS2017-MC
(b) NHTS2017-TP
(c) NHTS2017-CO
(d) LTDS2015-MC
(e) SGP2017-MC
Fig. 5.: Prediction accuracy by sample size, datasets, and outputs
Figures 5.3 and 5.3 demonstrate that predictive performance exhibits significant variation across datasets and shows a moderate increase with larger sample sizes. While model families display multi-modal performance distributions in Figure 5.2, the performance distributions of datasets appear more concentrated around the average in Figure 5.3. The highest accuracy is attained in LTDS2015 (76.58%), which is approximately 20 percentage points higher than NHTS2017 and SGP2017 data. The mean and median values of all three distributions are close, indicating a relatively concentrated distribution within each data set. Furthermore, predictive performance generally improves with sample size, although the rate of improvement appears to diminish slightly. For example, when the sample size is 1,000, the mean prediction accuracy is 51.63%, which increases to around 55.29% and 57.42% as the sample size grows to 10,000 and 100,000. This trend of marginally diminishing improvement holds consistent across datasets and outputs, as demonstrated in all five subfigures of Figure 5.3.
@@ 6 Limitations
This benchmark study has important limitations regarding data sources, experiment design, evaluation metrics, and research scope. While our literature and experiment data are relatively comprehensive, they are not exhaustive. The literature data excludes previous studies focusing solely on ML or DCM and those centered on spatiotemporal travel data [Mozolin2000, Polson2017, WuYuankai2018]. Our experiments exclude certain dimensions, such as training algorithms, hyper-parameters, feature selection, and feature transformation, which could influence model performance and change the benchmark results in our tournament model. Additionally, this empirical benchmark study uses prediction accuracy and model ranking as the evaluation criteria, neglecting other metrics like log-likelihood or Brier score [Abdar2021, Gruber2022]. Probabilistic metrics could be more appropriate for model evaluation from a theoretical perspective. However, since log-likelihood as a probabilistic metric does not exist in many deterministic ML models, prediction accuracy is the only metric that enables our comprehensive comparison of both deterministic and probabilistic models. The authors compared only DNNs and DCMs using log-likelihood and Brier scores in Appendix V, with a more comprehensive evaluation with probabilistic metrics left for future studies. Furthermore, model comparison should not solely focus on predictive performance. Recent studies have explored innovations in deep learning architectures [Sifringer2020, Han2022, Wong2021], investigated the potential of images and natural language [Cranenburgh2022, Wang2023], and examined interpretability and robustness as critical dimensions for model evaluation [Lipton2016, WangShenhao2020_econ_info, Wang2021_dnn_stat_learning]. Lastly, although we find the potential publication biases in favor of ML models’ higher performance by comparing literature and experiment data, our own experiment data are not devoid of publication biases. Publication bias appears an inherently challenging topic because any manuscript, including the unpublished ones, contains the authors’ intention of being published, and the existing tests for publication biases are only statistical heuristics, which are largely inapplicable to the ML models [Thornton2000]. Unfortunately, these limitations cannot be addressed in this empirical benchmark work and need be investigated by future researchers.
@@ 7 Conclusions
This empirical benchmark study improves upon the research frontier of the deterministic model comparisons through innovations in methodology, findings, and implications. Regarding methodology, the tournament model emulates and yet generalizes the past deterministic comparisons. It resembles the comparative studies because modelers also seek to identify the model of higher performance through pairwise comparison. Yet, this tournament model generalizes the deterministic comparison by expanding the sample size, examining the impacts of contextual complexity, and adopting formal statistical methods to investigate the effects of model families for both literature and experiment data points. From a modeler’s perspective, such a probabilistic approach captures the multiple sources of randomness in model comparison, including data collection, selection of training and testing data, stochastic algorithms, and inherent randomness in every predictive task [Dietterich1998]. In fact, statistical meta-modeling approaches have been widely adopted in the machine learning community for comparing classifiers across data sets, with methods including pairwise comparison [Bradley1952_rank_analysis], block-wise tests [Demsar2006_stat_comparison], or recently Bayesian approaches [Wainer2023_bayesian_comparison]. From a mathematical perspective, the past deterministic comparisons are only specific cases of our tournament models. The tournament model degenerates into a deterministic comparative study when the variance of $\epsilon_{m}$ approaches zero, the contextual factors are removed, and the sample size of pairwise comparisons is reduced to only one observation. In fact, if a deterministic process represents the underlying data generating process, our probabilistic tournament model should yield very high predictive power (e.g., 100% $\rho^{2}$ when no omitted variable exists). However, our empirical finding suggests otherwise, reaching a $\rho^{2}$ of only 0.4. Such an empirical result suggests that the performance comparison is much more random than a pure deterministic process.
@@ Acknowledgement
[…]
Appendix I: Summary of Acronyms and Terminology
Table 6: Acronyms and other terminology in this study;
Panel 1. Acronyms
DCM Discrete choice models
DNN Deep neural networks
DA Discriminant analysis
BM Bayesian models
SVM Support vector machines
KNN K nearest neighbors
DT Decision tree
GLM Generalized linear models
GP Gaussian process
BAGGING Bagging methods
RF Random forests
BOOSTING Boosting methods
MNL Multinomial logit
NL Nested logit
MXL Mixed logit
NHTS2017 National household travel survey 2017
LTDS2015 London travel demand survey 2015
SGP2017 Singapore stated preference survey 2017
MC Mode choice
CO Car ownership
TP Trip purposes
Panel 2. Other terminology
Classifiers Models using discrete values as outputs
Hyperparameters A parameter used to control the learning process
0-1 loss An indicator function that returns 1 when the target and output are not equal and zero otherwise
Appendix II: Further Implementation Details
In our experiments, the inputs were selected and normalized before modeling. In the NHTS2017 data, the most important 50 features were chosen from the initial 115 by using $\chi^{2}$ feature selection, which examines the dependence between inputs and outputs with condensed information. In the LTDS2015 data, the best 40 features were chosen from the initial 48 features using the same method. The SGP2017 data has only 22 independent variables, so it is unnecessary to limit the number of input features. After the feature selection, all the input variables were normalized to Gaussian distributions centered on zero with unit variance.
The number of experiments is smaller than the theoretically maximum, because of the limitations on data availability, data storage, and computational difficulty. Although the theoretical maximum of experiment points is $105\times 3\times 3\times 3=2,835$, our experiments only examined 1,394 experiment points because of computational difficulties. Specifically, trip purposes and car ownership are not available in the LTDS2015 and SGP2017 datasets. For the three GP classifiers, the complexity for storage space is $\mathcal{O}(N^{2})$, where $N$ is the sample size, so the GP classifiers with larger than 1K sample sizes can lead to memory errors in a personal computer (PC) with 32G RAM. The MXL model can only be implemented for the 1K sample size, because its training time can exceed 10 hours when the sample size reaches 10K. Five-fold cross-validation was used to compute the average prediction accuracy in the testing sets for each classifier, leading to a total number of $1,394\times 5=6,970$ experiments. For each experiment, the data was split into the training and testing sets with a ratio of $4:1$.
Appendix III: Linear Model Results
Linear regressions can also be used with prediction accuracy as outputs, and model and contextual factors as inputs. The results are summarized in Table 7. The results demonstrate similar conclusions from our tournament model. The ML models outperform DCMs in both literature and experiment data, although the literature typically have higher confidence in the outperformance than the experiment data. The contextual factors and inherent randomness still dominate, as shown by the statistical significance and the high explanatory power of the contextual factors, and the large residual randomness. Due to the high similarity to the tournament model, we skip detailed discussions here.
Table 7: Results from Linear Models
Parameters Linear Models
Literature (NC) Literature (C) Experiment (NC) Experiment (C)
Constant 0.701 (0.020) *** 0.915 (0.034) *** 0.545 (0.017) *** 0.165 (0.003) ***
Model family effects
BAGGING 0.205 (0.110) 0.267 (0.094) ** 0.010 (0.023) 0.006 (0.014)
BM 0.000 (0.066) 0.037 (0.054) -0.068 (0.020) *** -0.071 (0.012) ***
BOOSTING 0.139 (0.034) *** 0.124 (0.029) *** 0.026 (0.018) 0.022 (0.011) *
DA NA NA -0.004 (0.019) -0.005 (0.011)
DNN 0.063 (0.024) * 0.095 (0.021) *** 0.033 (0.018) 0.029 (0.011) **
DT 0.150 (0.030) *** 0.102 (0.026) *** 0.009 (0.018) 0.005 (0.011)
GLM -0.221 (0.110) * -0.022 (0.095) 0.008 (0.019) 0.004 (0.012)
GP NA NA -0.061 (0.031) -0.052 (0.019) **
KNN 0.069 (0.110) 0.100 (0.090) -0.061 (0.022) ** -0.065 (0.013) ***
RF 0.176 (0.037) *** 0.160 (0.032) *** 0.026 (0.026) 0.022 (0.016)
SVM 0.130 (0.031) *** 0.093 (0.026) *** -0.080 (0.020) *** -0.081 (0.012) ***
Contextual effects
If_Asian -0.046 (0.022) * 0.190 (0.004) ***
If_AM -0.185 (0.042) *** NA
If_Europe -0.126 (0.022) *** 0.296 (0.004) ***
If_CO -0.105 (0.035) ** 0.116 (0.004) ***
if_TP -0.088 (0.031) ** 0.084 (0.004) ***
Input dimention -0.004 (0.001) ** 0.004 (0.000) ***
Num of alternatives -0.019 (0.006) ** 0.012 (0.002) ***
Sample size -0.347 (1.690) 3.302 (0.407) ***
Statistical summary
Num of Observations 178 178 1388 1388
AIC -243.879 -315.761 -2391.3 -3762.7
$R^{2}$ 0.23 0.53 0.11 0.671
Adjusted $R^{2}$ 0.178 0.473 0.103 0.667
Appendix IV: Collecting and Describing Literature Data
The process of literature data collection could influence our results. To enhance the replicability of this study, the authors present the detailed data collection process as follows. We identified the relevant studies by combining the results from Google, Google Scholar, Web of Science, and Scopus, thus overcoming the limit of the searching scope in an individual engine. We combined the methodological and topical keywords in our searching: the methodological keywords included “prediction”, “machine learning”, “discrete choice models”, and “deep learning”, and the topical key words include “travel behavior”, “travel demand”, “choice analysis”, “travel mode choice”, amongst others. Although such searching often yields a large number of results, the authors had to filter out an overwhelmingly large number of irrelevant studies, which typically adopted spatiotemporal methods for aggregate demand modeling, different from the cross-sectional data sets for individual behavioral analysis [^5] In the literature searching, the authors did not iterate over different spellings, e.g., “behavior” and “behaviour”, thus potentially missing some relevant studies.. With the initial set of studies, we iterated through snowball sampling by searching the related studies in their sections of “Related Work”, “Literature Review”, “Bibliography”. Such iterations were repeated until no new relevant paper can be identified. The online search leads to 35 studies that compared the predictive performance between classical discrete choice models and machine learning classifiers.
The scope of the literature data is limited to only those studies that compared machine learning and discrete choice models for individual travel behaviors in a cross-sectional context. As a result, three groups of studies were excluded in the searching process above. The first group are those using complex data structures different from classical cross-sectional data, such as the spatial, temporal, or spatiotemporal data structures for urban mobility. The second group are those that tackle aggregate travel decisions rather than the disaggregate individual travel decisions. The third group uses either only machine learning for prediction or only discrete choice models for travel behavioral analysis without comparing their performance. In fact, the three groups constitute a vast volume of studies, which are beyond the scope of this empirical benchmark work.
Figures 7 and 7 visualize the histogram of the best models and the prediction accuracy of the models in the literature data. The results reinforce our findings that the ensemble methods and the DNNs have the highest predictive performance, while the DCMs rank at the medium or lower ends of the predictive distributions. As shown in Figure 7, 16 out of the 35 studies found that DNNs achieve the highest performance and 11 of them found that the ensemble methods (RF, BOOSTING, and BAGGING) outperform others. The ensemble methods achieve the highest prediction accuracy, and the accuracy of the three ensemble families can be as high as nearly 99.86%, as shown in Figure 7. DNNs’ average prediction accuracy is around 76.47%, which can be as high as 98.4%. The MNL and NL achieve about 72.49% and 68.53% on average, which are about 4-8 percentage points lower than the DNNs and about 15-20 percentage points lower than the ensemble methods. However, similar to the experiment data, the performance of the model families also exhibit quite significant randomness, as shown by the large variation of model performance in each model family. Overall, the descriptive analysis reveals similar patterns between the literature and experiment data.
Fig. 6.: Histogram of the best models
Fig. 7.: Prediction accuracy of the examined models
Appendix V: Comparing DCMs and DNNs using Log-likelihood and Brier Score
To complement the findings using only the deterministic metrics, we offer a further comparison using probabilistic metrics, including log-likelihood and Brier scores [^6] Brier score is computed as $\frac{1}{N}\sum_{i=1}^{N}(y_{i}-P_{i})^{2}$, for the DCMs and DNNs. Table 8 summarizes the average scores across all the prediction tasks and data sets, with the best model colored in gray. Due to the computational limitations for MXL, only the data sets with 1K and 10K observations are presented. Our findings here are consistent with those in the manuscript. The models with the highest performance are typically the DNNs, regardless of the sample size or the probabilistic metric. Although the DNNs on average tend to underperform the DCMs when the sample size is small, they outperform the DCMs when the sample size reaches 10K, implying the importance of a relatively large sample size for complex model families such as DNNs. The variation of performance in the DCM family is quite limited, while that in the DNN family is much larger. This finding is also consistent with that in Figure 5.2.
Table 8: Comparing DNNs and DCMs using Log-likelihood and Brier score
Log-likelihood Brier score
1k 10k 1k 10k
mnl_B -218.62 -1953.65 0.1211 0.1166
nl_B -223.72 -1998.27 0.1192 0.1129
mxl_B -226.58 N.A. 0.1245 N.A.
DCM (all) -223.03 -1978.44 0.1214 0.1145
DNN_30_1_P -208.02 -1975.75 0.1167 0.1116
DNN_30_3_P -226.88 -1895.53 0.1189 0.1073
DNN_30_5_P -259.98 -1946.90 0.1286 0.1096
DNN_100_1_P -217.87 -1947.88 0.1165 0.1104
DNN_100_3_P -256.64 -1840.05 0.1271 0.1046
DNN_100_5_P -356.07 -1843.43 0.1454 0.1048
DNN_200_1_P -216.54 -1839.00 0.1150 0.1111
DNN_200_3_P -264.94 -1822.61 0.1280 0.1037
DNN_200_5_P -391.92 -1835.39 0.1490 0.1045
DNN (all) -267.875 -1883.95 0.1276 0.1074
Appendix VI: Investigating Input Dimension and Performance
Here we investigate how input dimensions influence model performance. Figure 7 visualizes the relationship between prediction accuracy and input dimensions using all the literature data (Figure 7a) and those with input dimensions higher than five (Figure 7b). In Figure 7a, prediction accuracy decreases with input dimension, consistent with the tournament model in the manuscript. However, such a result could be influenced by the potential outliers. Therefore in Figure 7b, after removing the outliers with input dimensions lower than five, the relationship becomes flat, indicating only a weak correlation between accuracy and input dimensions. After running a regression, we obtain an estimated function as $y_{i}=\beta_{0}+\beta_{1}x_{i}$, in which $\beta_{1}=-0.0024$ with p value equals to 0.22, implying an insignificant statistical relationship between accuracy and input dimensions after the outliers are removed.
We further track the outlier articles to understand how such studies achieve very high performance using a limited number of features [Shafique2015, Tang2015]. It is because these studies investigate research questions different from the common discrete choice modeling that uses either stated or revealed preference data. For example, Shafique (2015) infers travel mode choice using observed speed data, thus leading to high predictive performance [Shafique2015]. However, since such studies predict travel behavior by comparing the ML and DCMs, they still fall into our research scope. Nonetheless, we acknowledge that the high heterogeneity and insufficient contextual controls in our two data sets can lead to unstable results, which shall be tackled by future studies.
(a) Full literature data
(b) Literature data (input dimension ¿ 5)
Fig. 8.: Input dimension and prediction accuracy for the literature data

### v3

@@ 5.3 Examining performance across data sets, outputs, and sample sizes
The substantial uncertainty within model families and individual models demonstrates the importance of investigating the contextual factors’ impacts on model performance. Therefore, we further decompose the predictive performance along the contextual dimensions. Figure 4 visualizes the distributions of prediction accuracy in three data sets, and Figure 5 examines the prediction accuracy regarding data sets, outputs, and sample sizes. Table 5 reports the model ranking and performance in three panels and nine sub-tables. Within each sub-table, we report the top-10 models with the DCMs highlighted below, conditioning on specific data sets, outputs, and sample sizes. With such analysis, we could examine the performance consistency of individual models and model families across contexts.
Figure 4: Prediction accuracy by data sets (dashed and dotted lines represent mean and median values)
Figure 5: Prediction accuracy by sample size, datasets, and outputs
Figures 4 and 5 demonstrate that predictive performance exhibits significant variation across datasets and shows a moderate increase with larger sample sizes. While model families display multi-modal performance distributions in Figure 2, the performance distributions of datasets appear more concentrated around the average in Figure 4. The highest accuracy is attained in LTDS2015 (76.58%), which is approximately 20 percentage points higher than NHTS2017 and SGP2017 data. The mean and median values of all three distributions are close, indicating a relatively concentrated distribution within each data set. Furthermore, predictive performance generally improves with sample size, although the rate of improvement appears to diminish slightly. For example, when the sample size is 1,000, the mean prediction accuracy is 51.63%, which increases to around 55.29% and 57.42% as the sample size grows to 10,000 and 100,000. This trend of marginally diminishing improvement holds consistent across datasets and outputs, as demonstrated in all five subfigures of Figure 5.
@@ 6 Limitations
This benchmark study has important limitations regarding data sources, experiment design, evaluation metrics, and research scope. While our literature and experiment data are relatively comprehensive, they are not exhaustive. The literature data excludes previous studies focusing solely on ML or DCM and those centered on spatiotemporal travel data [67, 68, 69]. Our experiments exclude certain dimensions, such as training algorithms, hyper-parameters, feature selection, and feature transformation, which could influence model performance and change the benchmark results in our tournament model. Additionally, this empirical benchmark study uses prediction accuracy and model ranking as the evaluation criteria, neglecting other metrics like log-likelihood or Brier score [70, 71]. Probabilistic metrics could be more appropriate for model evaluation from a theoretical perspective. However, since log-likelihood as a probabilistic metric does not exist in many deterministic ML models, prediction accuracy is the only metric that enables our comprehensive comparison of both deterministic and probabilistic models. The authors compared only DNNs and DCMs using log-likelihood and Brier scores in Appendix V, with a more comprehensive evaluation with probabilistic metrics left for future studies. Furthermore, model comparison should not solely focus on predictive performance. Recent studies have explored innovations in deep learning architectures [72, 73, 74], investigated the potential of images and natural language [75, 76], and examined interpretability and robustness as critical dimensions for model evaluation [77, 8, 78]. Lastly, although we find the potential publication biases in favor of ML models’ higher performance by comparing literature and experiment data, our own experiment data are not devoid of publication biases. Publication bias appears an inherently challenging topic because any manuscript, including the unpublished ones, contains the authors’ intention of being published, and the existing tests for publication biases are only statistical heuristics, which are largely inapplicable to the ML models [79]. Unfortunately, these limitations cannot be addressed in this empirical benchmark work and need be investigated by future researchers.
@@ 7 Conclusions
This empirical benchmark study improves upon the research frontier of the deterministic model comparisons through innovations in methodology, findings, and implications. Regarding methodology, the tournament model emulates and yet generalizes the past deterministic comparisons. It resembles the comparative studies because modelers also seek to identify the model of higher performance through pairwise comparison. Yet, this tournament model generalizes the deterministic comparison by expanding the sample size, examining the impacts of contextual complexity, and adopting formal statistical methods to investigate the effects of model families for both literature and experiment data points. From a modeler’s perspective, such a probabilistic approach captures the multiple sources of randomness in model comparison, including data collection, selection of training and testing data, stochastic algorithms, and inherent randomness in every predictive task [80]. In fact, statistical meta-modeling approaches have been widely adopted in the machine learning community for comparing classifiers across data sets, with methods including pairwise comparison [81], block-wise tests [82], or recently Bayesian approaches [83]. From a mathematical perspective, the past deterministic comparisons are only specific cases of our tournament models. The tournament model degenerates into a deterministic comparative study when the variance of $\epsilon_{m}$ approaches zero, the contextual factors are removed, and the sample size of pairwise comparisons is reduced to only one observation. In fact, if a deterministic process represents the underlying data generating process, our probabilistic tournament model should yield very high predictive power (e.g., 100% $\rho^{2}$ when no omitted variable exists). However, our empirical finding suggests otherwise, reaching a $\rho^{2}$ of only 0.4. Such an empirical result suggests that the performance comparison is much more random than a pure deterministic process.